the inclintion of a straight line is150 degree;if the perpendicular distance of the line from the origin is 10 unit,find the equation of the straight line.

Grade:11

1 Answers

Akshay
8 years ago
y=mX+c,
m=tan(150)=-tan(30)=-1/rt(3),
put(0,0) point, |0-0-c|/(rt(1+m2)=10,
c=+/-10rt(10)/3